OpenStudy (anonymous):

Please help whats the best way to solve this 4-s/9-s = 5/s-9

OpenStudy (anonymous):

is this \[\frac{4-s}{9-x}=\frac{5}{x-9}\]?

OpenStudy (anonymous):

if so start with \[\frac{s-4}{s-9}=\frac{5}{s-9}\] so \[s-4=5\] \[s=9\] but s cannot be 9 because that will make the denominator = 0 so there is no solution

OpenStudy (anonymous):

in other words it is not possible that \[\frac{4-s}{9-s}=\frac{5}{s-9}\]